# Quantum Mechanics Relative to a Quantum Reference System: a Relative State Approach

### v2

@@ I Introduction
Because the vast body of literature on the subject on the relational formulations and internal evolution of quantum mechanics, it is hardly for us to fully grasp all the existing works to review here, we only provide a potentially incomplete overview of some key landmark studies that we consider crucial and relevant to the topic of this paper, and how our works differ from these aforementioned studies are leaving in the section-VII. The first internal clock time evolution formalism of quantum mechanics can be attributed to the seminal work of Page and Wootters (PhysRevD.27.2885, ; 1986Quantum, ) in 1980s. And there are also a large subsequent literature on the Page-Wootters mechanism, relational time and quantum clock, see e.g. (Foti:2020erm, ; Woods2019pagewootters, ; qfns-48vq, ) and references therein. The concept of quantum reference frame is also a long-standing research in literature, including early foundational works (PhysRevD.30.368, ) and extensive developments on superselection rules (RevModPhys.79.555, ), relational observables (Rovelli:1990pi, ), frame transformations (Giacomini:2017zju, ), perspective dependence (2020Quant…4..225V, ; Hohn_2020, ) and symmetry-based formulations (Loveridge:2017pcv, ) and large number of relevant references therein. And the quantum reference frame relies on a relational interpretation of quantum physics, which started from Rovelli’s seminal paper (Rovelli:1995fv, ). On the other hand, through analogies between the Hilbert space and the Riemannian spaces endowed with additional geometric structures (such as complex and symplectic structures), numerous attempts have been made to geometrize quantum mechanics and incorporate it into the realm of pure geometry (see (Kibble:1978tm, ; Provost:1980nc, ; Ashtekar:1997ud, ; Brody:1999cw, ; Grabowski:2005my, ; Bengtsson:2006rfv, ) and the references therein). Among these, (Pandya:2006mrv, ; Matassa:2020dnc, ) place particular emphasis on the application of the Fubini-Study metric, a special metric in the Complex Projective Space, in describing quantum states, as well as the geometry underlying non-dynamical phases of quantum states (Simon:1983mh, ; Page:1987ca, ; Anandan:1990fq, ; Artacho:2017zey, ). Meanwhile, (Mosseri:2001gbm, ; Levay:2003pb, ; Grabowski:2005my, ) focus on understanding quantum entangled states from a geometric perspective. There have been many endeavors to establish an intrinsic, background-independent, and geometrized quantum mechanics. (Minic:2003en, ; Minic:2003nx, ) propose a quantum evolution equation akin to Einstein’s equation and provide a geometrized interpretation of probabilities. The problem of time in covariant quantum gravity is given by e.g. (Rovelli:1995fv, ; Gambini:2006ph, ), and the conditional probability interpretations of the Hamiltonian constraint is given by e.g. (Dolby:2004ak, ).
@@
In Section II of this paper, we first examine, from a historical perspective, how clocks, as a standard reference system, have been introduced and evolved in physics. In Section III, we discuss how to interpret the relative relationship between the quantum state of object’s coordinates and the quantum state of a clock from the perspective of the entangled states. We also explore the general relative interpretation of the entangled states and its underlying geometry: the non-trivial fiber bundles. In Section IV, from a purely geometric standpoint, we discuss how the quantum state of object’s coordinates evolves relative to the quantum state of the clock. In Section V, the relationship between the derived evolution equation and the Schrödinger equation is discussed. In Section VI, several non-inertial effects that arise in the reference frame of this quantum clock is discussed. We give some comparision between our work and the past works in Section VII. Finally, we conclude the paper with a summary. In the appendix, we review the functional integral approach, which also deals with quantum clocks, for the purpose of comparing it with the relative state approach.
@@ II Clock Time as a Reference System
Before going into details how a clock time measuring the so-called “time”, let us first take a broad look at the concept of spacetime as a reference system based on the quantum equivalence principle (2024arXiv240809630L, ). In fact, we have never “truly” measured spacetime itself, all our measurements of the so-called spacetime are indirect, relying on some material reference system (such as light or other material fields serving as references). A quantum equivalence principle (2024arXiv240809630L, ) provides the physical foundation for establishing a quantum reference frame theory through the measurement of the spacetime using material reference systems. Based on this principle, we know how to use material measuring rods and clocks to measure those universal properties (independent of the mass and other properties of the material reference system) that can be reinterpreted as the geometric properties of the spacetime, rather than merely properties of the measuring rods and clocks themselves. These universal properties of the spacetime reflected in material reference systems can manifest not only at the classical level as first-order moment averages independent of the material reference system’s mass (such as the averaged coordinates) but also at the quantum level as second-order moment quantum fluctuations of the material reference system (such as the coordinate broadening). Attributing these universal properties of material reference systems to spacetime may stem from human’s priori instinct or a desire for conceptual simplicity, akin to how, in early human history, barter trading (“trading goods for goods”) was inconvenient when comparing the values of different items. Subsequently, the development of a universal currency, which abstracts and measures the general value in all transactions and serves as a general equivalent, greatly simplified trading. Similarly, the spacetime reference frame plays the role of a general equivalent for comparing different motions. Clock time, as the simplest example of a spacetime reference frame, is the focus of this paper. After clarifying the example of the clock time, the concept of spacetime reference frames can naturally be extended in a similar manner.
@@ III Entangled State as a relative state
For simplicity, let’s consider a one-dimensional moving object with a spatial coordinate $X(\tau)$, whose motion with respect to the parameter $\tau$ is governed by a Hamiltonian $H_{X}$. Additionally, we now introduce a clock, with its pointer coordinate $T(\tau)$ evolving with respect to $\tau$ as described by a Hamiltonian $H_{T}$. Here, the global parameter $\tau$ shared by both $X$ and $T$ can be interpreted as a global Newtonian parameter. Furthermore, we assume that there is no interaction between the object $X(\tau)$ and the clock $T(\tau)$. They evolve independently with respect to the parameter $\tau$. Thus, the Hamiltonian of the entire system is given by $H=H_{X}+H_{T}$. The quantum state space of the entire system is the direct product space
@@
Only when the probability of the object’s position state $\textrm{P}\left(X(\tau)|T\right)$, equals the probability obtained from separate measurements of the object’s position and the clock, i.e., when the state is separable i.e. $C_{\tau}|^{2}=|A_{\tau}|^{2}|B_{\tau}|^{2}$, can we use a single global time to cover all possible position states. In this case, the entangled state $|X,T\rangle$ reverts to the separable direct product state $|X\rangle\otimes|T\rangle$ or a trivial fiber bundle, allowing the entire background state of the clock to be separated out and reduced to the standard textbook quantum state $|X\rangle$, which describes the probability $|B_{\tau}|^{2}$ of measuring the object’s position $X$ at a global time $\tau$ ($|B_{\tau}|^{2}$ has physical meaning w.r.t. $\tau$). Therefore, we require a relativistic quantum mechanical interpretation based on the “relative relationship between the two states $|X(\tau)\rangle$ and $|T(\tau)\rangle$” described by the local entangled states, to generalize the Copenhagen interpretation of quantum mechanics for “absolute quantum states $|X(\tau)\rangle$” in textbooks. This is analogous to generalizing global Cartesian geometry to curved Riemannian geometry based on tangent bundles over local base spaces. Establishing a relativistic quantum theory and an intrinsic interpretation that measures one entangled subspace (the quantum system under-study) relative to another subspace (the quantum measuring instrument) within a larger quantum state space is particularly urgent in explaining a quantum theory of gravity.
@@
$\langle T(\tau)|T(\tau^{\prime})\rangle=\langle T(\tau)|T\rangle\langle T|T(\tau^{\prime})\rangle=A^{\tau*}A^{\tau^{\prime}}=\frac{A^{\tau*}}{A_{\tau^{\prime}}}=A^{\tau}A^{\tau^{\prime}*}=\frac{A^{\tau}}{A_{\tau^{\prime}}^{*}}=s^{\tau\tau^{\prime}}=\left(s\right)_{\tau\tau^{\prime}}^{-1}\neq\delta_{\tau\tau^{\prime}}$ (5)
@@
$\langle T|X,T\rangle=\sum_{\tau}C_{\tau}|X(\tau)\rangle\langle T|T(\tau)\rangle=\sum_{\tau,\tau^{\prime}}C_{\tau}A_{\tau}^{*}\langle T(\tau^{\prime})|T(\tau)\rangle|X(\tau)\rangle=\sum_{\tau,\tau^{\prime}}C_{\tau}A_{\tau}^{*}s^{\tau\tau^{\prime}}|X(\tau)\rangle=\sum_{\tau}\frac{C_{\tau}}{A_{\tau}}|X(\tau)\rangle$ (8)
@@
Here, taking the partial trace of the entangled state’s density matrix is not simply summing the diagonal elements within a subspace, but rather contracting by the metric $s^{\tau\tau^{\prime}}$of the subspace $T$, i.e. $\textrm{Tr}_{T}\left(\mathbf{O}\right)=\mathbf{O}_{\mu\nu}s^{\mu\nu}\neq\mathbf{O}_{\mu\nu}\delta_{\mu\nu}$. This is analogous to how, in a curved Riemannian space, taking the trace of a matrix is no longer a straightforward summation of all diagonal elements, instead, it requires contraction using the curved metric. This represents a significant difference between quantum mechanics based on entangled states and conventional (global) quantum mechanics.
@@
The entangled state between the under-study system and that of a quantum instrument, along with its interpretation in terms of relative probabilities, provides the conceptual foundation for a quantum theory formulated in a general spacetime coordinate system. Entangled states are well-defined within the framework of quantum theory, yet they exhibit many properties that appear strange from the perspective of classical physics. In the early years of the quantum mechanics, it was assumed that the scale of the wave function was limited to atomic scale, and thus there was no need to worry about the issue of simultaneity arising from the instantaneous collapse of the wave function in the Copenhagen interpretation. However, Einstein’s EPR thought experiment pointed out that the scale of the wave function could, in fact, be much larger—so large that the wave functions are completely macroscopically orthogonal in space. This leads to the perplexing problem of “superluminal” effects associated with the instantaneous collapse of the wave function. Here, we merely offer an alternative perspective on the entangled state, viewing them as a type of “relative state” that describes the “relationship” between two states. This approach makes the picture of entangled state less incomprehensible. Take, for example, the spin version of the EPR thought experiment. A pair of spin-polarized photons or electrons, after undergoing a brief interaction—such as when light passes through a nonlinear medium, emitting two oppositely moving polarized photons A and B, leading to an entangled state, ofter called, the EPR state
@@ IV Relative Evolution of States and Kähler-Einstein Equation
It is worth noting that the prepared entangled state 12 is not a standard Schmidt decomposed entangled state. Here, $C_{i}$ is not Schmidt coefficient, and the local subspace basis vectors, which are selected from the calibration process, are not necessarily orthonormal. In contrast, the Schmidt decomposition requires selecting an appropriate local coordinate system $|X_{i}^{\prime}\rangle|T_{i}^{\prime}\rangle$ in the total quantum state space such that the expansion coefficients $C_{i}^{\prime}$ of the entangled state in this coordinate system are real-valued, and the local basis vectors are orthonormal.
@@
Due to the Hermiticity of the metric, the metric components vanish when both indices are unbarred or both indices are barred. The basis vectors that are locally dual to the local basis vectors $|X_{i}\rangle$ are now denoted as $\langle X^{j}|$, and hence the bar in the bra vector of previous eq.(6) can be omitted in this convention (one upper index and one lower index). They satisfy the orthogonality and normalization conditions
@@
Let’s examine the physical interpretation of the metric solutions of the equations. During the instrument calibration prior to measurement, mutually entangled quantum subsystems under-study $|X_{i}\rangle\in\mathscr{H}_{X}$ and the quantum instrument $|T_{a}\rangle\in\mathscr{H}_{T}$ are prepared. The entangled state they form can be written as 12. During the measurement process following instrument calibration, there is no further interaction between the quantum subsystem under-study and the quantum instrument. The coefficients $C_{i}$ and initial conditions for $h_{\bar{i}j}$ are established during the instrument calibration process and do not change with respect to any (hypothetical) external parameters afterward. The metric $h_{\bar{i}j}$ merely reflects the relative changes between the quantum subsystem under-study $X$ and the quantum instrument $T$. When the quantum instrument is in a certain quantum state $|T\rangle$ , the solutions to the equation provide us with the metric $h_{\bar{i}j}$ or basis vectors $|X_{i}\rangle$ of the subsystem under-study that is entangled with its basis vectors $|T_{a}\rangle$. Conversely, when the system is in a certain quantum eigenstate $|X\rangle$, the equation (28) yields the metric $s_{\bar{a}b}$ or basis vectors $|T_{a}\rangle$ that is entangled with its basis vectors $|X_{i}\rangle$. Thus, by utilizing the metric $s_{\bar{a}b}$ of eq.(28) or basis vectors $|T_{a}\rangle$, we can perform corresponding conditional projections on the entangled state $|X,T\rangle$ vector
@@
The diagonal elements of the matrix obtained after performing the partial trace directly represent the relative probabilities (or conditional probabilities) for the under-study subsystem in the state $|X_{i}\rangle$ when the quantum instrument is in the quantum state $|T\rangle$, yielding $\textrm{P}\left(X_{i}|T\right)=\left|\frac{C_{i}}{t_{i}}\right|^{2}$. As stressed before, the relative probability $\left|\frac{C_{i}}{t_{i}}\right|^{2}$ is not necessarily the standard absolute probability $|x_{i}|^{2}=h_{\bar{i}i}(t)$ being a solution of eq.(27). Only in the separable limit of the entangled state, it recovers the standard absolute probability $\textrm{P}\left(X_{i}|T\right)\rightarrow\textrm{P}\left(X_{i}\right)=|x_{i}|^{2}=h_{\bar{i}i}(t)$.
@@
$\textrm{Tr}_{T}\left[\rho_{X,T}\mathbf{O}(X)\right]=\sum_{i}\left|\frac{C_{i}}{t_{i}}\right|^{2}\langle X_{i}|\mathbf{O}|X_{i}\rangle$
@@
We observe that the relative probabilities naturally emerge during the process of taking partial traces over subsystems for the entangled state $|X,T\rangle$. This differs from the partial tracing of entangled states in Schmidt decomposition as presented in standard quantum mechanics. Firstly, for entangled states in standard Schmidt decomposition, partial tracing by the Kronecker’s delta, the two reduced density matrices obtained by taking partial traces over two distinct subsystems share the same non-zero eigenvalues. In contrast, in our case, taking partial traces over different subsystems by each non-trivial metric yields different relative probabilities. Secondly, the subsystem basis vectors of the entangled state in the Schmidt decomposition are orthonormal, whereas the subsystem basis vectors of our prepared entangled states are not orthonormal but instead define a non-trivial subspace metric. This also leads to the fact that the metric of our subsystems satisfies the general evolution equation 27 or 28 for local basis vectors, rather than being associated with the constant-curvature Fubini-Study metric (which satisfies the constant-curvature Einstein equation for complex projective spaces), as is the case for Schmidt’s entangled states.

### v3

@@ I Introduction
Because the vast body of literature on the subject on the relational formulations and internal evolution of quantum mechanics, it is hardly for us to fully grasp all the existing works to review here, we only provide a potentially incomplete overview of some key landmark studies that we consider crucial and relevant to the topic of this paper, and how our works differ from these studies are leaving in the section-VII. The first internal clock time evolution formalism of quantum mechanics can be attributed to the seminal work of Page and Wootters (PhysRevD.27.2885, ) in 1980s. And there are also a large subsequent literature on the Page-Wootters mechanism, relational time and quantum clock, see e.g. (Foti:2020erm, ; Woods2019pagewootters, ; qfns-48vq, ; Calcinari:2024pek, ) and references therein. The concept of quantum reference frame is also a long-standing research in literature, including early foundational works (PhysRevD.30.368, ) and extensive developments on superselection rules (RevModPhys.79.555, ), relational observables (Rovelli:1990pi, ), frame transformations (2013Changing, ; Giacomini:2017zju, ; Lake:2023nua, ; Carette:2023wpz, ; Castro-Ruiz:2025yvi, ), superposition of quantum frames (delaHamette:2021iwx, ), perspective dependence (2020Quant…4..225V, ; Hoehn:2019fsy, ; Hohn_2020, ) and symmetry-based formulations (Loveridge:2017pcv, ) and large number of relevant references therein. And the quantum reference frame relies on a relational interpretation of quantum physics, which started from Rovelli’s seminal paper (Rovelli:1995fv, ). On the other hand, through analogies between the Hilbert space and the Riemannian spaces endowed with additional geometric structures (such as complex and symplectic structures), numerous attempts have been made to geometrize quantum mechanics and incorporate it into the realm of pure geometry (see (Kibble:1978tm, ; Provost:1980nc, ; Ashtekar:1997ud, ; Brody:1999cw, ; Grabowski:2005my, ; Bengtsson:2006rfv, ) and the references therein). Among these, (Pandya:2006mrv, ; Matassa:2020dnc, ) place particular emphasis on the application of the Fubini-Study metric, a special metric in the Complex Projective Space, in describing quantum states, as well as the geometry underlying non-dynamical phases of quantum states (Simon:1983mh, ; Page:1987ca, ; Anandan:1990fq, ; Artacho:2017zey, ). Meanwhile, (Mosseri:2001gbm, ; Levay:2003pb, ; Grabowski:2005my, ) focus on understanding quantum entangled states from a geometric perspective. There have been many endeavors to establish an intrinsic, background-independent, and geometrized quantum mechanics. (Minic:2003en, ; Minic:2003nx, ) propose a quantum evolution equation akin to Einstein’s equation and provide a geometrized interpretation of probabilities. The problem of time in covariant quantum gravity is given by e.g. (Rovelli:1995fv, ; Gambini:2006ph, ), and the conditional probability interpretations of the Hamiltonian constraint is given by e.g. (Dolby:2004ak, ).
@@
In Section II of this paper, we first examine, from a historical perspective, how clocks, as a standard reference system, have been introduced and evolved in physics. In Section III, we discuss how to interpret the relative relationship between the quantum state of object’s coordinates and the quantum state of a clock from the perspective of the entangled states. We also explore the general relative interpretation of the entangled states and its underlying geometry: the non-trivial fiber bundles. In Section IV, from a purely geometric standpoint, we discuss how the quantum state of object’s coordinates evolves relative to the quantum state of the clock. In Section V, the relationship between the derived evolution equation and the Schrödinger equation is discussed. In Section VI, several non-inertial effects that arise in the reference frame of this quantum clock is discussed. We give some comparison between our work and the past works in Section VII. Finally, we conclude the paper with a summary. In the appendix, we review the functional integral approach, which also deals with quantum clocks, for the purpose of comparing it with the relative state approach.
@@ II Clock Time as a Reference System
Before going into details how a clock time measuring the so-called “time”, let us first take a broad look at the concept of spacetime as a reference system based on the quantum equivalence principle (2024arXiv240809630L, ). In fact, we have never “really” measured spacetime itself, all our measurements of the so-called spacetime are indirect, relying on some material reference system (such as light or other material fields serving as references). A quantum equivalence principle (2024arXiv240809630L, ) provides the physical foundation for establishing a quantum reference frame theory through the measurement of the spacetime using material reference systems. Based on this principle, we know how to use material measuring rods and clocks to measure those universal properties (independent of the mass and other properties of the material reference system) that can be reinterpreted as the geometric properties of the spacetime, rather than merely properties of the measuring rods and clocks themselves. These universal properties of the spacetime reflected in material reference systems can manifest not only at the classical level as first-order moment averages independent of the material reference system’s mass (such as the mean value of coordinate) but also at the quantum level as second-order moment quantum fluctuations of the material reference system (such as the broadening/variance of coordinate). Attributing these universal properties of material reference systems to spacetime may stem from human’s priori instinct or a desire for conceptual simplicity, akin to how, in early human history, barter trading (“trading goods for goods”) was inconvenient when comparing the values of different items. Subsequently, the development of a universal currency, which abstracts and measures the general value in all transactions and serves as a general equivalent, greatly simplified trading. Similarly, the spacetime reference frame plays the role of a general equivalent for comparing different motions. Clock time, as the simplest example of a spacetime reference frame, is the focus of this paper. After clarifying the example of the clock time, the concept of spacetime reference frames can naturally be extended in a similar manner.
@@ III Entangled State as a relative state
For simplicity, let us consider a one-dimensional moving object with a spatial coordinate $X(\tau)$, whose motion with respect to the parameter $\tau$ is governed by a Hamiltonian $H_{X}$. Additionally, we now introduce a clock, with its pointer coordinate $T(\tau)$ evolving with respect to $\tau$ as described by a Hamiltonian $H_{T}$. Here, the global parameter $\tau$ shared by both $X$ and $T$ can be interpreted as a global Newtonian parameter. Furthermore, we assume that there is no interaction between the object $X(\tau)$ and the clock $T(\tau)$. They evolve independently with respect to the parameter $\tau$. Thus, the Hamiltonian of the entire system is given by $H=H_{X}+H_{T}$. The quantum state space of the entire system is a direct product space
@@
Only when the probability of the object’s position state $\textrm{P}\left(X(\tau)|T\right)$, equals the probability obtained from separate measurements of the object’s position and the clock, i.e., when the state is separable i.e. $|C_{\tau}|^{2}=|A_{\tau}|^{2}|B_{\tau}|^{2}$, can we use a single global time to cover all possible position states. In this case, the entangled state $|X,T\rangle$ reverts to the separable direct product state $|X\rangle\otimes|T\rangle$ or a trivial fiber bundle, allowing the entire background state of the clock to be separated out and reduced to the standard textbook quantum state $|X\rangle$, which describes the probability $|B_{\tau}|^{2}$ of measuring the object’s position $X$ at a global time $\tau$ ($|B_{\tau}|^{2}$ has physical meaning w.r.t. $\tau$). Therefore, we require a relativistic quantum mechanical interpretation based on the “relative relationship between the two states $|X(\tau)\rangle$ and $|T(\tau)\rangle$” described by the local entangled states, to generalize the Copenhagen interpretation of quantum mechanics for “absolute quantum states $|X(\tau)\rangle$” in textbooks. This is analogous to generalizing global Cartesian geometry to curved Riemannian geometry based on tangent bundles over local base spaces. Establishing a relativistic quantum theory and an intrinsic interpretation that measures one entangled subspace (the quantum system under-study) relative to another subspace (the quantum measuring instrument) within a larger quantum state space is particularly urgent in explaining a quantum theory of gravity.
@@
$\langle T(\tau)|T(\tau^{\prime})\rangle=\langle T(\tau)|T\rangle\langle T|T(\tau^{\prime})\rangle=A^{\tau*}A^{\tau^{\prime}}=\frac{A^{\tau*}}{A_{\tau^{\prime}}}=A^{\tau}A^{\tau^{\prime}*}=\frac{A^{\tau}}{A_{\tau^{\prime}}^{*}}=s^{\tau\tau^{\prime}}=\left(s_{\tau\tau^{\prime}}\right)^{-1}\neq\delta_{\tau\tau^{\prime}}$ (5)
@@
$\langle T|X,T\rangle=\sum_{\tau}C_{\tau}|X(\tau)\rangle\langle T|T(\tau)\rangle=\sum_{\tau,\tau^{\prime}}C_{\tau}A_{\tau^{\prime}}^{*}\langle T(\tau^{\prime})|T(\tau)\rangle|X(\tau)\rangle=\sum_{\tau,\tau^{\prime}}C_{\tau}A_{\tau^{\prime}}^{*}s^{\tau\tau^{\prime}}|X(\tau)\rangle=\sum_{\tau}\frac{C_{\tau}}{A_{\tau}}|X(\tau)\rangle$ (8)
@@
Here, taking the partial trace of the entangled state’s density matrix is not simply summing the diagonal elements within a subspace, but rather contracting by the metric $s^{\tau\tau^{\prime}}$of the subspace $T$, i.e. $\textrm{Tr}_{T}\left(\mathbf{O}\right)=\mathbf{O}_{\mu\nu}s^{\mu\nu}\neq\mathbf{O}_{\mu\nu}\delta_{\mu\nu}$. This is analogous to how, in a curved Riemannian space, taking the trace of a matrix is no longer a straightforward summation of all diagonal elements, instead, it requires contraction using the curved metric. This represents a significant difference between this approach and the conventional (global) quantum mechanics.
@@
The entangled state between the under-study system and that of a quantum instrument, along with its interpretation in terms of the relative probabilities, provides the conceptual foundation for a quantum theory formulated in a general spacetime coordinate system. Entangled states are well-defined within the framework of quantum theory, yet they exhibit many properties that appear strange from the perspective of classical physics. In the early years of the quantum mechanics, it was assumed that the scale of the wave function was limited to atomic scale, and thus there was no need to worry about the issue of simultaneity arising from the instantaneous collapse of the wave function in the Copenhagen interpretation. However, Einstein’s EPR thought experiment pointed out that the scale of the wave function could, in fact, be much larger—so large that the wave functions are completely macroscopically orthogonal in space. This leads to the perplexing problem of “superluminal” effects associated with the instantaneous collapse of the wave function. Here, we merely offer an alternative perspective on the entangled state, viewing them as a type of “relative state” that describes the “relationship” between two states. This approach makes the picture of entangled state less incomprehensible. Take, for example, the spin version of the EPR thought experiment. A pair of spin-polarized photons or electrons, after undergoing a brief interaction—such as when light passes through a nonlinear medium, emitting two oppositely moving polarized photons A and B, leading to an entangled state, often called, the EPR state
@@ IV Relative Evolution of States and Kähler-Einstein Equation
It is worth noting that the prepared entangled state (12) is not a standard Schmidt decomposed entangled state. Here, $C_{i}$ is not Schmidt coefficient, and the local subspace basis vectors, which are selected from the calibration process, are not necessarily orthonormal. In contrast, the Schmidt decomposition requires selecting an appropriate local coordinate system $|X_{i}^{\prime}\rangle|T_{i}^{\prime}\rangle$ in the total quantum state space such that the expansion coefficients $C_{i}^{\prime}$ of the entangled state in this coordinate system are real-valued, and the local basis vectors are orthonormal.
@@
Due to the Hermiticity of the metric, the metric components vanish when both indices are unbarred or both indices are barred. The basis vectors that are locally dual to the local basis vectors $|X_{i}\rangle$ are now denoted as $\langle X^{j}|$, so in this notation, the bar over the bra vector $\bar{\langle X|}$ (dual to $|X\rangle$), as previous eq.(6), can be omitted in this convention. Throughout the following paper, the upper index and lower index dual to each other, as the notations we often used in Riemannian geometry, they satisfy the orthogonality and normalization conditions
@@
Let us examine the physical interpretation of the metric solutions of the equations. During the instrument calibration prior to measurement, mutually entangled quantum subsystems under-study $|X_{i}\rangle\in\mathscr{H}_{X}$ and the quantum instrument $|T_{a}\rangle\in\mathscr{H}_{T}$ are prepared. The entangled state they form can be written as (12). During the measurement process following instrument calibration, there is no further interaction between the quantum subsystem under-study and the quantum instrument. The coefficients $C_{i}$ and initial conditions for $h_{\bar{i}j}$ are established during the instrument calibration process and do not change with respect to any (hypothetical) external parameters afterward. The metric $h_{\bar{i}j}$ merely reflects the relative changes between the quantum subsystem under-study $X$ and the quantum instrument $T$. When the quantum instrument is in a certain quantum state $|T\rangle$ , the solutions to the equation provide us with the metric $h_{\bar{i}j}$ or basis vectors $|X_{i}\rangle$ of the subsystem under-study that is entangled with its basis vectors $|T_{a}\rangle$. Conversely, when the system is in a certain quantum eigenstate $|X\rangle$, the equation (28) yields the metric $s_{\bar{a}b}$ or basis vectors $|T_{a}\rangle$ that is entangled with its basis vectors $|X_{i}\rangle$. Thus, by utilizing the metric $s_{\bar{a}b}$ of eq.(28) or basis vectors $|T_{a}\rangle$, we can perform corresponding conditional projections on the entangled state $|X,T\rangle$ vector
@@
The diagonal elements of the matrix obtained after performing the partial trace directly represent the relative probabilities (or conditional probabilities) for the under-study subsystem in the state $|X_{i}\rangle$ when the quantum instrument is in the quantum state $|T\rangle$, yielding $\textrm{P}\left(X_{i}|T\right)=\left|\frac{C_{i}}{t_{i}}\right|^{2}$. As stressed before, the relative probability $\left|\frac{C_{i}}{t_{i}}\right|^{2}$ does not strictly equal to the standard absolute probability $\textrm{P}\left(X_{i}\right)=|x_{i}|^{2}=h_{\bar{i}i}(t)$ being a solution of eq.(27). Only in the separable limit of the entangled state (12) (or equivalently, the quantum clock can be seen as a global external inertial clock), it recovers the standard absolute probability $\textrm{P}\left(X_{i}|T\right)\overset{separable}{\approx}\textrm{P}\left(X_{i}\right)=\left\||X_{i}\rangle\right\|^{2}=h_{\bar{i}i}(t)$, and hence, only in this sense, solving the evolution equation (27) for $|X_{i}\rangle$ or $h_{\bar{i}j}$ is approximately equivalent to solve a Schrödinger-like equation for the probability of the particle’s position $\textrm{P}\left(X_{i}|T\right)\overset{separable}{\approx}\textrm{P}\left(X_{i}\right)$ (we will see the eq.(27) recovering a Schrödinger-like equation in the next section). But in strict sense, the evolution of the relative probability $\textrm{P}\left(X_{i}|T\right)$ must be given by the evolution of $|T_{a}\rangle$ (or $s_{\bar{a}b}$) w.r.t. $x$ according to eq.(28).
@@
$\textrm{Tr}_{T}\left[\rho_{X,T}\mathbf{O}(X)\right]=\sum_{i}\left|\frac{C_{i}}{t_{i}}\right|^{2}\langle X_{i}|\mathbf{O}|X_{i}\rangle$ (31)
@@
We observe that the relative probabilities naturally emerge during the process of taking partial traces over subsystems for the entangled state $|X,T\rangle$. This differs from the partial tracing of entangled states in Schmidt decomposition as presented in standard quantum mechanics. Firstly, for entangled states in standard Schmidt decomposition, partial tracing by the Kronecker’s delta, the two reduced density matrices obtained by taking partial traces over two distinct subsystems share the same non-zero eigenvalues. In contrast, in our case, taking partial traces over different subsystems by each non-trivial metric yields different relative probabilities. Secondly, the subsystem basis vectors of the entangled state in the Schmidt decomposition are orthonormal, whereas the subsystem basis vectors of our prepared entangled states are not orthonormal but instead define a non-trivial subspace metric. This also leads to the fact that the metric of our subsystems satisfies the general evolution equation (27) or (28) for local basis vectors, rather than being associated with the constant-curvature Fubini-Study metric (which satisfies the constant-curvature Einstein equation for complex projective spaces), as is the case for Schmidt’s entangled states.